Tyler has 2 saving accounts. The 2 accounts pay 10% and 12% in annual interest; there is $500.00 more in the account that pays 12% than there is in the other account. If the total interest for a year is $214, how much money does he have in each account?
Thank you.

Answer by scott8148(6628)

(10%) (x) + (12%) (x + 500) = 214 ___ multiply by 100 to clear percents

10x + 12x + 6000 = 21400

22x = 15400 ___ x = 700 ___ x + 500 = 1200